# Mitigating photon loss in linear optical quantum circuits: classical postprocessing methods outperforming postselection

### v2

@@ 2 Overview of main results
Our final contribution is in providing strong evidence that photon loss error mitigation techniques based on ZNE [32] cannot outperform postselection. Indeed, in section 11 we provide an upper bound on the error incurred from ZNE, $E_{extrap}$, which is larger than the worst-case statistical error of postselection (see Thm. 13). In appendix 21, we also provide numerical evidence that for all $n\geq n_{0}$, for some $n_{0}\in\mathbb{N}$, $E_{extrap}$ is always larger than the statistical error of postselection. Our result highlights the subtleties involved in adapting existing error mitigation techniques to mitigate photon loss errors in DVLOQC. Indeed, in DVLOQC there is a natural way to mitigate photon loss errors, postselection, which is not present for other types of errors in other types of hardware [32]. This therefore sets a benchmark that a photon loss mitigation technique in DVLOQC must satisfy to be useful, namely that it must outperform postselection. Our results indicate that recycling mitigation succeeds in mitigating beyond what postselection can offer whereas ZNE does not.
3 Comparison to recent work
Techniques based on ZNE have been applied to mitigate photon loss in Gaussian boson sampling experiments [15]. Interestingly, we find that ZNE-based techniques do not seem to provide an advantage over postselection when adapted to the DVLOQC setting. An important distinction between ZNE-based methods and recycling mitigation, which is presented in this work, is that the former have an associated sampling overhead while recycling mitigation does not. In ZNE, many data points are collected, each corresponding to the expectation value of an observable $\mathsf{O}$ computed from a noisy circuit at a given noise level. These points are used to compute, via extrapolation, the noiseless expectation value of $\mathsf{O}$. Computing each of these data points to a desired precision has a sample cost, while recycling mitigation uses information from the lossy output statistics that would otherwise be discarded. This means that recycling mitigation may be applied with no additional sampling overhead relative to postselection.
Finally, recent work [55] has provided a method for mitigating photon loss in the continuous variable (CV) setting, adapting pre-existing methods of probabilistic error cancellation [13, 36, 37, 38] to the CV setting. Although the techniques of [55] are applicable to the DVLOQC setting, these are solely for expectation value mitigation (weak mitigation) [56]. Our methods, by contrast, can perform both strong (full probability distribution mitigation) as well as weak mitigation, with a classical memory cost scaling with the size of the set of probabilities to be mitigated. Furthermore, we present analytical as well as extensive numerical evidence that our methods provably outperform postselection, whereas, to our knowledge, no comparison between the performance of the developed mitigation methods and postselection is presented in [55].
@@ 4 Preliminaries
(a)                                                       (b)
Figure 1: A schematic illustrating the main steps of the recycling mitigation protocol. (a) An input state |$\ket{\psi_{in}}$ of $n$ photons is introduced to a lossy $m$-mode linear optical interferometer implementing a unitary transformation. The classical measurement outcomes of each of the output modes is denoted by the classical $m-$bit string $s^{n}_{l}$. (b) The classical data set generated by repeatedly sampling from the quantum circuit is then used as input for classical postprocessing. This classical postprocessing consists of three stages. First, the data set is used to generate lossy probability estimators. These estimators are then used to construct recycled probability estimators, which are in turn then used to generate mitigated values.
@@
  • We assume that sampling occurs in the no-collision regime, where at most one photon occupies any output mode. This is approximately true for $m\geq O(n^{2})$ [6]. In this regime the samples are given by $\mathbf{S}=(s_{1},\dots,s_{m})$, with $s_{i}\in\{0,1\}$, and are therefore bit strings of length $m$.
The uniform loss model is a widely used error model for simulating photon loss, and is the standard assumption used when constructing loss-tolerant quantum error correcting codes [59]. It is also often assumed when deriving efficient classical algorithms for simulating lossy linear optical setups [60]. Working in the no-collision regime is primarily interesting for two reasons. Firstly, one can prove statements of quantum advantage in boson sampling in this regime [6]. And secondly, the fact that at most one photon occupies each mode allows us to assume the use of the standard and widely available threshold detectors, rather than number resolving ones which are currently challenging to practically implement. As a final note, while the no-collision assumption is a useful one, it is not necessary for our techniques to work. Indeed, as discussed in later parts of this paper, our techniques can be generalised to the case where more than one photon can occupy a mode.
@@
As the the size of the system is increased, the probability of postselecting non-lossy outcomes decays exponentially towards zero. One motivation for the error mitigation techniques we introduce, and which we will detail in subsequent sections, is that in the high loss regime, where $\eta>0.5$, lossy outputs with fewer than $n$ photons are considerably more likely than $n$–photon outputs. The errors $\epsilon_{stat}(s_{i}^{n-k})$ for $k\neq 0$ are therefore smaller in general than $\epsilon_{stat}(s_{i}^{n})$, by standard arguments from statistics such as Hoeffding and Chebyshev inequalities [26]. Consequently, the estimated probabilities for lossy $n-k$–photon outputs are more converged and have lower statistical error than those of (lossless) $n$–photon outputs. Furthermore, these lossy probabilities can contain information on the probabilities of the $n$–photon outputs.
@@
Recycling mitigation uses the $n-k$–photon probability estimates $\{\tilde{p}(s_{i}^{n-k})\}$, potentially for a range of $k$ values, and construct from these a mitigated $n$–photon probability distribution $\{p_{mit}(s_{j}^{n})\}$. To have any utility, a photon loss mitigation technique needs to outperform computing the $n$–photon probability estimates $\{\tilde{p}(s_{j}^{n})\}$ from the samples $N_{tot,0}$, which we will henceforth refer to as postselection on $n$–photon outputs, or just postselection. We will therefore use postselection as the benchmark to evaluate the performance of recycling mitigation. Postselection is, to our knowledge, the only technique being used to mitigate the effects of photon loss on current DVLOQC hardware [18]. Another factor motivating recycling mitigation is that it does not increase the overall sample cost relative to postselection. This contrasts favourably with many error mitigation results that have an accompanying sampling overhead [32].
5 Recycled probabilities
@@
The recycled probabilities are constructed from $n-k$ output photon statistics, where $k\in\{1,\ldots,n-1\}$. To explicitly analyse the signal of the ideal probability within the recycled probability, the recycled probabilities may be decomposed into a combination of an ideal $n$–photon output probability, and an interference term consisting of a mixture of other $n$–photon output probabilities from the distribution. We first describe the construction of the recycled probabilities from $n-k$ output photon statistics, which generalises for all $k$. We then describe the analytical decomposition of the recycled probabilities into $n$–photon output probabilities.
5.1 Construction of recycled probabilities from lossy outputs
@@
We now detail the construction of recycled probabilities from $n-k$–photon output statistics - that is, the output statistics in which exactly $k$ of $n$ photons have been lost. This construction should be applied to obtain the recycled probability distribution in an experiment. The recycled probability for bit string $s^{n}_{l}$ computed from $n-k$–photon output statistics is denoted $p_{R}^{k}(s^{n}_{l})$, with $k\in\{1,\ldots,n-1\}$. Hence there are $n-1$ recycled probabilities one can construct from lossy output statistics for any $n$–photon output bit string, one for each possible value of $k$. Performing the construction involves summing over the $n-k$ output photon bit string probabilities that relate to a particular ideal probability. Informally, this relation is that these are the probabilities of $n-k$–photon output bit strings that the ideal output bit string can be mapped to through the loss of $k$ photons. We now provide a formal statement of this relation.
@@
We first define a mapping procedure from each $n$–photon output bit string to a set of $n-k$–photon output bit strings. Each mapped set of bit strings represents the set of all possible states that the associated $n$–photon output state could become after losing $k$ photons. Let $\mathcal{M}_{\text{unocc.},k,i}$ be the subset of $\{1,\dots,m\}$ corresponding to the unoccupied modes of the output bit string $s^{n-k}_{i}$. That is, the set of indices of the modes $j\in\{1,\dots,m\}$ of the bit string for which $s_{j}=0$. The number of unoccupied modes for $n-k$–photon outputs is $|\mathcal{M}_{\text{unocc.},k,i}|=m-n+k$. Let $\overline{\mathcal{M}_{\text{unocc.},k,i}}$ be the complement of $\mathcal{M}_{\text{unocc.},k,i}$ in $\{1,\dots,m\}$, so that $\overline{\mathcal{M}_{\text{unocc.},k,i}}\cup\mathcal{M}_{\text{unocc.},k,i}=\{1,\dots,m\}$. The subset $\overline{\mathcal{M}_{\text{unocc.},k,i}}$ denotes the occupied modes of $s^{n-k}_{i}$, consisting of the set of indices of modes $j\in\{1,\dots,m\}$ for which $s_{j}=1$. As the bit string $s^{n-k}_{i}$ represents an $n-k$–photon output the number of occupied modes is $|\overline{\mathcal{M}_{\text{unocc.},k,i}}|=n-k$. We define the set $\mathcal{L}(s^{n}_{i}):=\{s^{n-k}_{j}|s^{n}_{i}\Rightarrow s^{n-k}_{j}\}$. And the set of all size $k$ subsets of $\overline{\mathcal{M}_{\text{unocc.},0,i}}$ is $\overline{\mathcal{S}_{0,i}}:=\{X\subset\overline{\mathcal{M}_{\text{unocc.},0,i}}|\hskip 3.00003pt|X|=k\}.$ The symbol $\Rightarrow$’ denotes the operation where, for every size $k$ subset $\{\overline{l_{1}},\ldots,\overline{l_{k}}\}\in\overline{\mathcal{S}_{0,i}}$, the $n$–photon output bit string $s_{i}^{n}$ is mapped to a new $n-k$–photon output bit string $s^{n-k}_{j}$ by replacing $s_{\overline{l_{i}}}=1$ with $s_{\overline{l_{i}}}=0$. In this case, the number of $k-$subsets of $\overline{\mathcal{M}_{\text{unocc.},0,i}}$ is (${n\choose k}$, and so $|\mathcal{S}_{0,i}|={n\choose k}$ and the size of the generated set of bit strings is $|\mathcal{L}(s^{n}_{i})|={n\choose k}$.
@@
The recycled probability for the bit string $s_{l}^{n}$ can be defined as the sum of probabilities of the (${n\choose k}$ bit strings $s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})$,
$\begin{split}p_{R}^{k}(s^{n}_{l})&:=\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i}).\\ \end{split}$ (3)
@@
$\begin{split}p_{R}^{k}(s^{n}_{l})&=\frac{1}{{m-n+k\choose k}}\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i}).\\ \end{split}$ (4)
@@
To illustrate how this construction might work in practice we now provide a small example. If we would like to compute the recycled probability for the bit string: 111000, in an experiment in which there are $m=6$ modes, $n=3$ input photons and the construction is being performed for $k=1$ lost photons. The recycled probability may be computed directly from eqn. 4 as being
@@
In practice, rather than using the set of exact probabilities, $\{p(s_{i}^{n-k})\}$, to compute recycled probabilities in the manner shown in eqn. 4, instead empirical estimates of the exact probabilities, $\{\tilde{p}(s_{i}^{n-k})\}$, are used. These are calculated from the set of measured experimental output bit strings, as described for eqn. 1, and so include statistical errors due to finite samples. This statistical error is an important consideration when making comparisons with postselection, and is later included in the analysis of the protocol performance. We now provide pseudocode detailing how to construct the recycled probability for a specific $n$ output photon bit string and value of $k$ from a set of $N$ output bit strings sampled from a given circuit.
input : The $n$–photon output bit string $s^{n}_{l}$ for which the recycled probability estimator is to be constructed, a set of $N$ output sample bit strings from the DVLOQC circuit $\{s_{j}\}_{j\in\{1,\ldots,N\}}$, and the choice of $k$ value indicating that $n-k$ output photon statistics be used for the construction.
1  Initialise variable $\tilde{p}_{R}^{k}(s^{n}_{l})\leftarrow 0$ for the recycled probability estimator to be computed.
2 Create a new set of bit strings by discarding all output bit strings from set $\{s_{j}\}_{j\in\{1,\ldots,N\}}$ except those for which the number of measured output photons was $n-k$, with the new list denoted $\{s_{l}\}_{l\in\{1,\ldots,N_{est,k}\}}$ where $N_{est,k}\leq N$.
3 Generate the set of $n-k$–photon lossy output bit strings $\mathcal{L}(s^{n}_{l})$.
4 for l = $1$ to $N_{est,k}$ do
5       Initialise variable $X_{s_{l}}\leftarrow 0$.
6       if $s_{l}\in\mathcal{L}(s^{n}_{l})$ then
7             $X_{s_{l}}\leftarrow X_{s_{l}}+1$.
8       end if
9
10 end for
11Update recycled probability estimator variable as $\tilde{p}_{R}^{k}(s^{n}_{l})\leftarrow\frac{X_{s_{l}}}{{m-n+k\choose k}N_{est,k}}$.
@@
5.2 Decomposition of recycled probabilities into $n$–photon output bit string probabilities
In later sections, the recycled probabilities are analysed in terms of their decomposition into $n$–photon output probabilities. For a given output bit string $s^{n}_{l}$, this representation allows explicit treatment of the signal of the ideal $n$–photon output probability $p(s^{n}_{l})$ within the recycled probability $p^{k}_{R}(s^{n}_{l})$. To get the recycled probabilities in this form, the lossy output probabilities within the sum in eqn. 4 are decomposed into $n$–photon output probabilities from the ideal distribution. The details of this decomposition will now be formalised.
For this purpose we now define another mapping procedure, this time from each lossy $n-k$–photon output bit string from the sum in eqn. 4 to a set of $n$–photon output bit strings. Where each mapped set of bit strings represents the set of $n$–photon outputs that the associated lossy output bit string could have been had loss not occurred. The set of all size $k$ subsets of $\mathcal{M}_{\text{unocc.},k,i}$ is defined $\mathcal{S}_{k,i}:=\{X\subset\mathcal{M}_{\text{unocc.},k,i}|\hskip 3.00003pt|X|=k\}$. We define the set $\mathcal{G}(s^{n-k}_{i}):=\{s_{j}^{n}|s_{i}^{n-k}\to s^{n}_{j}\}$. The symbol $\to$’ denotes the operation where, for every size $k$ subset $\{l_{1},\ldots,l_{k}\}\in\mathcal{S}_{k,i}$, the bit string $s_{i}^{n-k}$ is mapped to a new bit string by replacing each bit $s_{l_{i}}=0$ with $s_{l_{i}}=1$. The number of $k-$subsets of $\mathcal{M}_{\text{unocc.},k,i}$ is (${m-n+k\choose k}$, and so $|\mathcal{S}_{k,i}|={m-n+k\choose k}$ and the size of the generated set of bit strings is $|\mathcal{G}(s^{n-k}_{i})|={m-n+k\choose k}$.

### v3

@@ 2 Overview of main results
Our final contribution is in providing strong evidence that photon loss error mitigation techniques based on ZNE [17] cannot outperform postselection in the DVLOQC setting. Indeed, in Section 7 we provide an upper bound on the error incurred from ZNE, $E_{extrap}$, which is larger than the worst-case statistical error of postselection (see Thm. 16). In Appendix 15, we also provide numerical evidence that for all $n\geq n_{0}$, for some $n_{0}\in\mathbb{N}$, $E_{extrap}$ is always larger than the statistical error of postselection. Indeed, in DVLOQC there is a natural way to mitigate photon loss errors, postselection, which is not present for other types of errors in other types of hardware [17]. This therefore sets a benchmark that a photon loss mitigation technique in DVLOQC must satisfy to be useful, namely that it must outperform postselection.
@@ 3 Preliminaries
  • We assume that sampling occurs in the no-collision regime, where at most one photon occupies any output mode. This is approximately true for $m\in\Omega(n^{2})$ [1]. In this regime the samples are given by $\mathbf{S}=(s_{1},\dots,s_{m})$, with $s_{i}\in\{0,1\}$, and are therefore bit strings of length $m$. The total number of possible no-collision outputs is $m\choose n$}.
The uniform loss model is a widely used error model for simulating photon loss, and is the standard assumption used when constructing loss-tolerant quantum error correcting codes [60]. It is also often assumed when deriving efficient classical algorithms for simulating lossy linear optical setups [19]. Working in the no-collision regime is primarily interesting for two reasons. Firstly, one can prove statements of quantum advantage in boson sampling in this regime [1]. And secondly, the fact that at most one photon occupies each mode allows us to assume the use of the standard and widely available threshold detectors, rather than number resolving ones which are currently challenging to practically implement. As a final note, while the no-collision assumption is a useful one, it is not necessary for our techniques to work. Indeed, as discussed in later parts of this paper, our techniques can in principle be generalised to the case where more than one photon can occupy a mode.
@@
In postselection, estimators are constructed using only the outputs for which $k=0$, when photon loss is uniform this corresponds to approximately $N_{tot}(1-\eta)^{n}$ samples. As the the size of the system is increased, the probability of postselecting non-lossy outcomes decays exponentially towards zero.
For values of loss above a certain threshold, the statistical error of probabilities constructed from lossy statistics is in general lower than those constructed from lossless ones. As an example of how to see this, note that for $k=1$, the expected number of samples is $N_{tot}n(1-\eta)^{n-1}\eta$. For the lossless $k=0$ case, we have $N_{tot}(1-\eta)^{n}$ such samples. Since the statistical error is typically upper bounded by $1/\sqrt{N}$, where $N$ is the sample number, we can see that the inequality $N_{tot}n(1-\eta)^{n-1}\eta\geq N_{tot}(1-\eta)^{n}$ (which implies lower statistical errors for lossy probabilities) holds whenever $\eta\geq\frac{1}{n+1}$.
@@
Recycling mitigation uses the nkn-k–photon probability estimates $\{\tilde{p}(s_{i}^{n-k})\}$, potentially for a range of $k$ values, and construct from these a mitigated nn–photon probability distribution $\{p_{mit}(s_{j}^{n})\}$. To have any utility, a photon loss mitigation technique needs to outperform computing the nn–photon probability estimates $\{\tilde{p}(s_{j}^{n})\}$ from the samples $N_{tot,0}$, which we will henceforth refer to as postselection on nn–photon outputs, or just postselection. We will therefore use postselection as the benchmark to evaluate the performance of recycling mitigation. Postselection is, to our knowledge, the only technique being used to mitigate the effects of photon loss on current DVLOQC hardware [40]. Another factor motivating recycling mitigation is that it does not increase the overall sample cost relative to postselection. This contrasts favourably with many error mitigation results that have an accompanying sampling overhead [17].
4 Method
4.1 Recycled probabilities
@@
The recycled probabilities are constructed from nkn-k output photon statistics, where $k\in\{1,\ldots,n-1\}$. To explicitly analyse the signal of the ideal probability within the recycled probability, the recycled probabilities may be decomposed into a combination of an ideal nn–photon output probability, and an interference term consisting of a mixture of other nn–photon output probabilities from the distribution. We first describe the construction of the recycled probabilities from nkn-k output photon statistics, which generalises for all $k$. We then describe the analytical decomposition of the recycled probabilities into nn–photon output probabilities.
4.1.1 Construction of recycled probabilities from lossy outputs
@@
We now detail the construction of recycled probabilities from nkn-k–photon output statistics - that is, the output statistics in which exactly $k$ of $n$ photons have been lost. This construction should be applied to obtain the recycled probability distribution in an experiment. The recycled probability for bit string $s^{n}_{l}$ computed from nkn-k–photon output statistics is denoted $p_{R}^{k}(s^{n}_{l})$, with $k\in\{1,\ldots,n-1\}$. Hence there are n1n-1 recycled probabilities one can construct from lossy output statistics for any nn–photon output bit string, one for each possible value of $k$. Performing the construction involves summing over the nkn-k output photon bit string probabilities that relate to a particular ideal probability. Informally, this relation is that these are the probabilities of nkn-k–photon output bit strings that the ideal output bit string can be mapped to through the loss of $k$ photons. We now provide a formal statement of this relation.
@@
We first define a mapping procedure from each nn–photon output bit string to a set of nkn-k–photon output bit strings. Each mapped set of bit strings represents the set of all possible states that the associated nn–photon output state could become after losing $k$ photons. Let $\mathcal{M}_{\text{unocc.},k,i}$ be the subset of $\{1,\dots,m\}$ corresponding to the unoccupied modes of the output bit string $s^{n-k}_{i}$. That is, the set of indices of the modes $j\in\{1,\dots,m\}$ of the bit string for which $s_{j}=0$. The number of unoccupied modes for nkn-k–photon outputs is $|\mathcal{M}_{\text{unocc.},k,i}|=m-n+k$. Let $\overline{\mathcal{M}_{\text{unocc.},k,i}}$ be the complement of $\mathcal{M}_{\text{unocc.},k,i}$ in $\{1,\dots,m\}$, so that $\overline{\mathcal{M}_{\text{unocc.},k,i}}\cup\mathcal{M}_{\text{unocc.},k,i}=\{1,\dots,m\}$. The subset $\overline{\mathcal{M}_{\text{unocc.},k,i}}$ denotes the occupied modes of $s^{n-k}_{i}$, consisting of the set of indices of modes $j\in\{1,\dots,m\}$ for which $s_{j}=1$. As the bit string $s^{n-k}_{i}$ represents an nkn-k–photon output the number of occupied modes is $|\overline{\mathcal{M}_{\text{unocc.},k,i}}|=n-k$. We define the set $\mathcal{L}(s^{n}_{i}):=\{s^{n-k}_{j}|s^{n}_{i}\Rightarrow s^{n-k}_{j}\}$. And the set of all size $k$ subsets of $\overline{\mathcal{M}_{\text{unocc.},0,i}}$ is $\overline{\mathcal{S}_{0,i}}:=\{X\subset\overline{\mathcal{M}_{\text{unocc.},0,i}}|\hskip 3.00003pt|X|=k\}.$ The symbol $\Rightarrow$’ denotes the operation where, for every size $k$ subset $\{\overline{l_{1}},\ldots,\overline{l_{k}}\}\in\overline{\mathcal{S}_{0,i}}$, the nn–photon output bit string $s_{i}^{n}$ is mapped to a new nkn-k–photon output bit string $s^{n-k}_{j}$ by replacing $s_{\overline{l_{i}}}=1$ with $s_{\overline{l_{i}}}=0$. In this case, the number of $k$-subsets of $\overline{\mathcal{M}_{\text{unocc.},0,i}}$ is $n\choose k$}, and so $|\mathcal{S}_{0,i}|={n\choose k}$ and the size of the generated set of bit strings is $|\mathcal{L}(s^{n}_{i})|={n\choose k}$.
@@
The recycled probability for the bit string $s_{l}^{n}$ can be defined as the sum of probabilities of the $n\choose k$} bit strings $s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})$,
[…]
@@
[…]
@@
To illustrate how this construction might work in practice we now provide a small example. If we would like to compute the recycled probability for the bit string: 111000, in an experiment in which there are $m=6$ modes, $n=3$ input photons and the construction is being performed for $k=1$ lost photons. The recycled probability may be computed directly from eqn. 10 as being
@@
In practice, rather than using the set of exact probabilities, $\{p(s_{i}^{n-k})\}$, to compute recycled probabilities in the manner shown in eqn. 10, instead empirical estimates of the exact probabilities, $\{\tilde{p}(s_{i}^{n-k})\}$, are used. These are calculated from the set of measured experimental output bit strings, as described for eqn. 7, and so include statistical errors due to finite samples. This statistical error is an important consideration when making comparisons with postselection, and is later included in the analysis of the protocol performance. We now provide pseudocode detailing how to construct the recycled probability for a specific $n$ output photon bit string and value of $k$ from a set of $N$ output bit strings sampled from a given circuit.
input : The nn–photon output bit string $s^{n}_{l}$ for which the recycled probability estimator is to be constructed, a set of $N$ output sample bit strings from the DVLOQC circuit $\{s_{j}\}_{j\in\{1,\ldots,N\}}$, and the choice of $k$ value indicating that nkn-k output photon statistics be used for the construction.
1ex Initialise variable $\tilde{p}_{R}^{k}(s^{n}_{l})\leftarrow 0$ for the recycled probability estimator to be computed.
Create a new set of bit strings by discarding all output bit strings from set $\{s_{j}\}_{j\in\{1,\ldots,N\}}$ except those for which the number of measured output photons was nkn-k, with the new list denoted $\{s_{l}\}_{l\in\{1,\ldots,N_{est,k}\}}$ where $N_{est,k}\leq N$.
Generate the set of nkn-k–photon lossy output bit strings $\mathcal{L}(s^{n}_{l})$.
Initialise variable $X_{s_{l}}\leftarrow 0$.
for l = 11 to $N_{est,k}$ do
 if $s_{l}\in\mathcal{L}(s^{n}_{l})$ then
    $X_{s_{l}}\leftarrow X_{s_{l}}+1$.
   end if
end for
Update recycled probability estimator variable as $\tilde{p}_{R}^{k}(s^{n}_{l})\leftarrow\frac{X_{s_{l}}}{{m-n+k\choose k}N_{est,k}}$.
@@
4.1.2 Decomposition of recycled probabilities into nn–photon output bit string probabilities
In later sections, the recycled probabilities are analysed in terms of their decomposition into nn–photon output probabilities. For a given output bit string $s^{n}_{l}$, this representation allows explicit treatment of the signal of the ideal nn–photon output probability $p(s^{n}_{l})$ within the recycled probability $p^{k}_{R}(s^{n}_{l})$. To get the recycled probabilities in this form, the lossy output probabilities within the sum in eqn. 10 are decomposed into nn–photon output probabilities from the ideal distribution. The details of this decomposition will now be formalised.
For this purpose we now define another mapping procedure, this time from each lossy nkn-k–photon output bit string from the sum in eqn. 10 to a set of nn–photon output bit strings. Where each mapped set of bit strings represents the set of nn–photon outputs that the associated lossy output bit string could have been had loss not occurred. The set of all size $k$ subsets of $\mathcal{M}_{\text{unocc.},k,i}$ is defined $\mathcal{S}_{k,i}:=\{X\subset\mathcal{M}_{\text{unocc.},k,i}|\hskip 3.00003pt|X|=k\}$. We define the set $\mathcal{G}(s^{n-k}_{i}):=\{s_{j}^{n}|s_{i}^{n-k}\to s^{n}_{j}\}$. The symbol $\to$’ denotes the operation where, for every size $k$ subset $\{l_{1},\ldots,l_{k}\}\in\mathcal{S}_{k,i}$, the bit string $s_{i}^{n-k}$ is mapped to a new bit string by replacing each bit $s_{l_{i}}=0$ with $s_{l_{i}}=1$. The number of $k$-subsets of $\mathcal{M}_{\text{unocc.},k,i}$ is $m-n+k\choose k$}, and so $|\mathcal{S}_{k,i}|={m-n+k\choose k}$ and the size of the generated set of bit strings is $|\mathcal{G}(s^{n-k}_{i})|={m-n+k\choose k}$.